# A Reduced Basis Decomposition Approach to Efficient Data Collection in Pairwise Comparison Studies

### v1

@@ 5.2 Scheduling for classroom based peer learning
Comparative judgement is perhaps most used in educational context, as a form of peer learning and assessment. Under the theory of “Learning by Evaluation”, students carry out an exercise and then compare pairs of answers Bartholomew2020. The act of comparing peer work allows students to implicitly internalize quality standards and construct a robust understanding of the assessment criteria, often more effectively than through passive rubric analysis Hoffelinck2025.
@@
We demonstrate the utility of our proposed method using data from JonesSirl2017, where undergraduate mathematics students assessed peer responses to calculus conceptual problems. The data set consists of 3,258 comparisons of 139 students’ written responses to a first year calculus assignment. The comparisons were made by the first year student who undertook the assignment as part of a Learning by Evaluation study, with the aim of improving both their understanding of calculus and ability to write formal mathematical arguments. There are around 23 comparisons for each assignment.
We fit the Bradley-Terry model to the first 700 comparisons (approximately 20% of the total number of comparisons), using $\lambda_{i}\sim N(0,5^{2}),\,i=1,\ldots,139$ as the prior distributions on the quality of the assignments. We consider this as the first phase of the class, which is followed by an in-lecture activity or discussion. To run the second phase efficiently, we place a joint prior distribution on the quality parameters $\boldsymbol{\lambda}\sim N(\boldsymbol{0},C)$, where $C$ is the posterior covariance matrix from the first phase. We can then construct the approximate scheduling distribution $\widetilde{\mathcal{S}}$ and show the students pairs of assignments according to this distribution. Reading in the data and fitting the Bradley–Terry model takes 10.6 seconds. Using the standard method, it takes 15 minutes 43 seconds to construct $B$ and then compute $\mathcal{S}$, for a total workflow of 15 minutes 54 seconds. Using our RBD method, it takes 4.08 seconds to construct $\widetilde{\mathcal{S}}$ directly from $C$, for a total workflow time of 14.68 seconds and reducing the time taken by 98%.
@@ 6 Discussion
A limitation of our approach is that it assumes a static design framework with known prior information. While this accommodates sequential multi-phase designs where interim posteriors inform subsequent scheduling, it does not address fully online adaptive designs that update after every comparison. However, as noted in Section 1, static designs avoid the reliability inflation and increasing comparison difficulty problems associated with adaptive comparative judgement (Bramley2018; Jones2023). Another consideration is numerical precision: the sensitivity analysis (Section 4.3) demonstrates that very small tolerance values ($\varepsilon_{R}<10^{-12}$) can lead to precision errors in standard numerical computing environments, though the method remains robust across practically relevant tolerance ranges. Finally, regarding computational complexity, if $\operatorname{rank}(\boldsymbol{C})\ll N$, full precision for the few non-zero eigenvalues of $\Delta$ still requires N1N-1 iterations of the RBD method. Using fewer iterations will yield approximated spectral results rather than the exact values.
A broader limitation concerns the optimality of the underlying design criterion itself. Although D-optimal designs can now be computed in a classical framework (Roettger2025), we are not aware of any method to contract Bayesian D-optimal designs. The computational efficiency of our RBD method may allow for Bayesian D-optimal designs to be constructed for the first time. The principal-component-based design of Seymour25 assigns higher scheduling probability to pairs exhibiting greater prior variance, but it remains an open question whether this criterion is optimal in any formal sense, such as D-optimality or A-optimality. Establishing optimality would require comparing alternative design criteria, yet such comparisons have been hindered by the computational expense of constructing designs under the standard approach. By reducing this computational burden, our RBD method opens the door to systematic investigation of design optimality for comparative judgement studies. Researchers can now feasibly construct and evaluate competing design criteria for moderately large studies, enabling future work to establish theoretical foundations for optimal experimental design in the Bradley-Terry framework with correlated priors.
@@
7 Acknowledgements
This work was supported by a UKRI Future Leaders Fellowship [MR/X034992/1] and an EPSRC Mathematical Sciences Discipline Hopping Grant [UKRI2389]. We thank Cat Smith for her help in creating the coding workflow for this project. We thank Ian Jones for providing the Learning by Evaluation data.

### v2

@@ 5.2 Scheduling for classroom based peer learning
Comparative judgement is perhaps most used in educational contexts, as a form of peer learning and assessment. Under the theory of “Learning by Evaluation”, students carry out an exercise and then compare pairs of answers (Bartholomew et al., 2020). The act of comparing peer work allows students to implicitly internalize quality standards and construct a robust understanding of the assessment criteria, often more effectively than through passive rubric analysis (Hoffelinck et al., 2025).
@@
We demonstrate the utility of our proposed method using data from Jones and Sirl (2017), where undergraduate mathematics students assessed peer responses to calculus conceptual problems. The data set consists of 3,258 comparisons of 139 students’ written responses to a first year calculus assignment. The comparisons were made by the first year student who undertook the assignment as part of a Learning by Evaluation study, with the aim of improving both their understanding of calculus and ability to write formal mathematical arguments. There are around 23 comparisons for each assignment.
We fit the Bradley-Terry model to the first 700 comparisons (approximately 20% of the total number of comparisons), using $\lambda_{i}\sim N(0,5^{2}),\,i=1,\ldots,139$ as the prior distributions on the quality of the assignments. We consider this as the first phase of the class, which is followed by an in-lecture activity or discussion. To run the second phase efficiently, we place a joint prior distribution on the quality parameters $\boldsymbol{\lambda}\sim N(\boldsymbol{0},C)$, where $C$ is the posterior covariance matrix from the first phase. We can then construct the approximate scheduling distribution $\widetilde{\mathcal{S}}$ and show the students pairs of assignments according to this distribution. Reading in the data and fitting the Bradley–Terry model takes 10.6 seconds. Using the standard method, it takes 15 minutes 43 seconds to construct $B$ and then compute $\mathcal{S}$, for a total workflow of 15 minutes 54 seconds. Using our RBD method, it takes 4.08 seconds to construct $\widetilde{\mathcal{S}}$ directly from $C$, for a total workflow time of 14.68 seconds, thus reducing the time taken by 98%.
@@ 6 Discussion
A potential benefit of static scheduling designs in peer learning contexts, is that unlike most other comparative judgement scenarios, we care not only about the statistical efficiency of the design but also about the pedagogical experience of the student acting as judge. Comparisons between very similar pieces of work may be statistically informative but offer limited learning value, as the student has little to reflect on when the differences are negligible. Conversely, comparisons between work of vastly different quality may be too obvious to prompt meaningful engagement with the assessment criteria. A well-constructed scheduling distribution can balance these considerations.
@@
A limitation of our approach is that it assumes a static design framework with known prior information. While this accommodates sequential multi-phase designs where interim posteriors inform subsequent scheduling, it does not address fully online adaptive designs that update after every comparison. However, as noted in Section 1, static designs avoid the reliability inflation and increasing comparison difficulty problems associated with adaptive comparative judgement (Bramley and Vitello, 2018; Jones and Davies, 2023). Another consideration is numerical precision: the sensitivity analysis (Section 4.3) demonstrates that very small tolerance values ($\varepsilon_{R}<10^{-12}$) can lead to precision errors in standard numerical computing environments, though the method remains robust across practically relevant tolerance ranges. Finally, regarding computational complexity, if $\operatorname{rank}(\boldsymbol{C})\ll N$, full precision for the few non-zero eigenvalues of $\Delta$ still requires N1N-1 iterations of the RBD method. Using fewer iterations will yield approximated spectral results rather than the exact values.
A broader limitation concerns the optimality of the underlying design criterion itself. Although D-optimal designs can now be computed in a classical framework (Röttger et al., 2025), we are not aware of any method to contract Bayesian D-optimal designs. The computational efficiency of our RBD method may allow for Bayesian D-optimal designs to be constructed for the first time. The principal-component-based design of Seymour et al. (2025) assigns higher scheduling probability to pairs exhibiting greater prior variance, but it remains an open question whether this criterion is optimal in any formal sense, such as D-optimality or A-optimality. Establishing optimality would require comparing alternative design criteria, yet such comparisons have been hindered by the computational expense of constructing designs under the standard approach. By reducing this computational burden, our RBD method opens the door to systematic investigation of design optimality for comparative judgement studies. Researchers can now feasibly construct and evaluate competing design criteria for moderately large studies, enabling future work to establish theoretical foundations for optimal experimental design in the Bradley-Terry framework with correlated priors.
@@
Acknowledgements
This work was supported by a UKRI Future Leaders Fellowship [MR/X034992/1] and an EPSRC Mathematical Sciences Discipline Hopping Grant [UKRI2389]. We thank Cat Smith for her help in creating the coding workflow for this project. We thank Ian Jones for providing the Learning by Evaluation data. We would like to thank the reviewers for their insightful comments and constructive feedback, which have improved the quality of this work.